How can I solve the following 2 questions:
1.Find the 2 consecutive integers equal to 77.
2. 5-2t = M

Answer by longjonsilver(2297)

let first number be x, so next number must be x+1.

So, x+(x+1) = 77
2x+1 = 77
2x = 76
x = 76/2
x = 38.

So 2 numbers are 38 and 39

2. You have 2 unknowns, t and M with just one equation. You cannot solve this. You need more info. Which variable are you trying to find? t or M?

Jon.